Circular No. 6415

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 T. Dotani, H. Negoro, K. Mitsuda, H. Inoue, and F. Nagase, Institute

of Space and Astronautical Science, communicate on behalf of the ASCA team: “We observed Cyg X-1 with ASCA during May 30.23–31.14 UT with a net exposure time of ∼ 33 000 s. Cyg X-1 was in a high, soft state during the observations. The GIS intensity in the range 1–10 keV was ∼ 1000 counts/s (after deadtime correction), which corresponds to ∼ 1.3 Crab. The observed flux in 1–10 keV was 3.2 × 10^{−8} erg s^{−1} cm^{−2}. The energy spectrum can be approximated by a blackbody with a hard tail. The bestfitting temperature of the blackbody is 0.34 ± 0.02 keV, and the hard tail can be approximated by a power law with a photon index of 2.4 ± 0.1. The blackbody and the power-law components cross over around 2.3 keV, above which the power-law component becomes dominant.”

GRO J1744–28 J. M. Kommers, R. E. Rutledge, D. W. Fox, W. H. G. Lewin, and

E. H. Morgan, Massachusetts Institute of Technology; C. Kouveliotou, Universities Space Research Association; and J. van Paradijs, University of Alabama in Huntsville and University of Amsterdam, report: “We found strong quasiperiodic oscillations (QPOs) with a frequency of ∼ 0.4 Hz in a burst from GRO J1744–28 observed on ∼ June 4.617 UT with the PCA instrument onboard the Rossi X-ray Timing Explorer. The burst was the sixth of seven seen on June 4. The oscillations are clearly visible in the lightcurve during the last 25 s of the 40-s burst. A power spectrum of these 25 s shows a peak near 0.4 Hz, with FWHM < 0.08 Hz (the upper limit being set by our frequency resolution) and a fractional rms amplitude of the mean burst flux during these 25 s (in excess of the quiescent flux) of 25 ± 5 percent (not corrected for deadtime). We have also identified QPOs with frequencies in the range 0.3–0.5 Hz in the emission following at least 10 of 82 bursts observed between late January and late April. These QPOs typically lasted for ∼ 30 s, although in one burst they lasted as long as 80 s; the fractional rms amplitudes exceeded 15 percent in some cases.”

COMET C/1996 B2 (HYAKUTAKE) Naked-eye m_{1} estimates by D. Seargent, The Entrance, N.S.W.: May

17.81 UT, 3.9; 21.80, 4.1; 23.81, 4.3; 26.79, 4.4; 28.79, 4.5; 30.79, 4.8.

1996 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green